A motor boat traveling with a current went 35 miles in 3.5 hours. Traveling agaist the current, the boat went 12 miles in 3 hours. Let r be the rate of the boat un calm water and let w be the rate of the current. Write both your solutions and equations you used to solve the problem.
Answer by tommyt3rd(5050)

1) (r+w)(3.5)=35 so r+w=10
2) (r-w)(3)=12, so r-w=4

adding equations...
2r=14
r=7mph

so w=3mph